Problem 6: A heat exchanger is to be used in a heating process. A standard type of heat exchanger (code name HX1) has an installed cost of $20,000 and a useful life of 6 years. At the end of its useful life, the standard heat exchanger has a negligible salvage value, assumed to be 0. Another type of heat exchanger (Code name HX2) with equivalent design capacity is priced at S34,000 ihstalled, has a useful life of 10 years and a salvage value of $4,000. Assuming an effective interest rate of 6% per year, compounded discretely and that the replacement cost of each exchanger is the same as that of the original exchanger, determine which heat exchanger is the lower cost based capitalized cost analysis.
Answer #1

First determine the LCM of 6 and 10. It is equal to 30 years.

Now determine the PW of both the machines

PW(HX1) = - 20,000-20,000(P/F,6%,6) -20,000(P/F,6%,12)- 20,000(P/F,6%,18)- 20,000(P/F,6%,24)

PW (HX1) = -$ 55,985

Similarly for machine 2

PW(HX2) = -34,000 -30,000(P/F,6%,10) - 30,000(P/F,6%,20) + 4000(P/F,6%,30)

PW (HX2,) =- $ 59,409.54

Capitalized cost = PW/I

Machine HX1, Capitalized cost = -$55,985/0.06 = -$ 933,083

machine HX2, capitalized cost = -$59409.54/0.06 = - $ 990,159

Select machine HX1.